OpenStudy (anonymous):

Jerry plans to begin saving money for a trip by putting $1 in a savings account the first month and then adding enough to double the amount in the account each following month. a. After he finishes contributing to the account, he will withdraw $500 to make a down payment on the trip. Create a function to show how much money will be in the account if he finishes contributing to the account after t months.

OpenStudy (gnorris):

The amount he deposits is a geometric sequence: 1, 2, 4, 8, 16, ... This is a geometric sequence with a first term of 1 and a common ratio of 2. a = 1 r = 2 The general formula for the sum of the first t terms of a geometric series is: S(t) = a(1 - r^t) / (1 - r) Plug in your values for a and r: S(t) = 1(1 - 2^t) / (1 - 2) S(t) = (1 - 2^t) / -1 S(t) = -1 + 2^t S(t) = 2^t - 1 Answer: S(t) = 2^t - 1 P.S. If he did this for 9 months, he'd have saved $511. P.P.S. If he did this for 2 years, he'd have $16,777,215. P.P.P.S. If he did this for 3 years, he'd have the net worth of Bill Gates at around $70 billion. P.P.P.P.S. If he did this for 4 years, he'd have about $281 trillion. That's enough to give each person in the U.S. about $1 million each.